# TopoBench: A Framework for Benchmarking Topological Deep Learning

### v2

@@ 1 Introduction
In geometric deep learning (GDL; Bronstein et al., 2021), graph neural networks (GNNs; Zhou et al., 2020) have demonstrated impressive capabilities in processing relational data represented as graphs. However, because graphs represent relationships through edges, they inherently capture only pairwise interactions, which can be a limiting factor. For example, social interactions often involve groups of individuals rather than just pairs, and electrostatic interactions in proteins can span multiple atoms. Topological deep learning (TDL; Papamarkou et al., 2024; Bodnar, 2023; Hajij et al., 2023; Papillon et al., 2023) offers a framework for modeling complex systems characterized by such multi-way relations among components, leveraging to that end higher-order discrete topological domains (such as simplicial and cell complexes, see Section 2). Topological neural networks (TNNs; Feng et al., 2019; Bunch et al., 2020; Hajij et al., 2020; Bodnar et al., 2021a; Ebli et al., 2020; Schaub et al., 2021; Bodnar et al., 2021b; Chien et al., 2021), which are part of TDL, have found applications in numerous fields that involve higher-order relational data such as social networks (Knoke and Yang, 2019), protein biology (Jha et al., 2022), physics (Wei and Fink, 2024), and computer networks (Bernárdez et al., 2025). TNNs have also shown their potential in various machine learning tasks (Dong et al., 2020; Barbarossa and Sardellitti, 2020; Chen et al., 2022; Roddenberry et al., 2021; Telyatnikov et al., 2023; Giusti et al., 2023).
@@ 2 Background
Figure 2: Topological Deep Learning Domains. Nodes in blue, (hyper)edges in pink, and faces in dark red. Figure adopted from Papillon et al. (2023).
@@ 3 Existing Software
Various tools also exist for higher-order domains. For hypergraphs, simplicial complexes, and other topological structures, HyperNetX (Liu et al., 2021), XGI (Landry et al., 2023), DHG (Feng et al., 2019), and TopoX (Hajij et al., 2024) each focus on different facets. HyperNetX facilitates hypergraph computations, whereas XGI supports both hypergraphs and simplicial complexes. DHG implements deep learning algorithms for graphs and hypergraphs. TopoX is a suite of three packages—TopoNetX, TopoEmbedX, and TopoModelX—that provide broader support for hypergraphs, simplicial, cellular, path, and combinatorial complexes. TopoNetX facilitates constructing and computing on these domains, including working with nodes, edges, and higher-order cells; TopoEmbedX embeds higher-order domains into Euclidean spaces, while TopoModelX implements most TNNs surveyed in Papillon et al. (2023).
@@ 4 The TopoBench Library: Module Outline, Datasets and Liftings
TopoBench implements a unified and flexible workflow that facilitates the addition of new datasets, data manipulation and preprocessing methods (collectively referred to as transforms), and deep learning models. This design ensures applicability across a wide range of tasks and enables a broad cross-domain comparison, currently lacking in the TDL literature. Each module within TopoBench is assigned a distinct role while maintaining a consistent input-output structure, which provides a modular interface across all topological domains. Figure 1 outlines the TopoBench modules, grouped by functionality into data, model, training, and communication components. Algorithm 1 illustrates the TopoBench execution pipeline in pseudo-code.
@@
6:      nn.Encoder(cfg.model),
7:      nn.Backbone(cfg.model),
8:      nn.BackboneWrapper(cfg.model),
9:      nn.Readout(cfg.model),
@@ 4.1 TopoBench Modules
The readout module converts latent representations from the neural network into final predictions. The Loss module defines a loss function (from the PyTorch library, or customized), while the optimizer module configures the optimizer and scheduler. This design allows seamless use of any optimizer and scheduler from torch.optim, thereby supporting flexible and robust training. Finally, the evaluator module, built upon torchmetrics, provides metrics for both classification and regression tasks –while also allowing for tailored ones for specific datasets and tasks.
@@ 4.2 Datasets
Datasets with higher-order interactions. Several datasets with higher-order interactions are included in TopoBench, showcasing its capabilities to handle data supported on hypergraphs, simplicial complexes, and other topological domains. The Mantra dataset (Ballester et al., 2024) is part of TopoBench, offering over 43,000 two-dimensional and 249,000 three-dimensional triangulations of surfaces and manifolds, which can be used, for example, as features on a simplicial complex. In addition, the widely used AllSet hypergraph datasets (Chien et al., 2021)—Cora-Cocitation, Citeseer-Cocitation, PubMed-Cocitation, Cora-Coauthorship, and DBLP-Coauthorship—are integrated, following the same preprocessing as HyperGCN (Yadati et al., 2019). These hypergraphs group documents co-authored or co-cited together into single hyperedges. Collectively, these examples illustrate how TopoBench supports data beyond traditional graph pairwise interactions.
@@ 4.3 Topological Liftings
Table LABEL:tab:submissions in Appendix B provides a comprehensive list of all of the liftings currently implemented in TopoBench. Currently, TopoBench supports 11 structural liftings targeting simplicial complexes; 2 targeting cell complexes; 10 moving to hypergraphs; and 3 from a domain to combinatorial complexes. Notably, TopoBench’s modular design simplifies the integration of additional liftings, ensuring the framework’s adaptability to evolving research needs.
@@ 5 Numerical Experiments
This section presents numerical experiments that illustrate the breadth of TopoBench’s functionality by performing a cross-domain comparison. [^3] All code and instructions to reproduce the experiments of this section are publicly available; see Appendix Code Availability and Reproducibility for details on code availability and reproducibility. The overall setup is first described, then results from benchmarking various graph, hypergraph, and TNNs are reported, and an ablation study on signal propagation is presented to demonstrate how TopoBench supports comparisons in TDL.
@@ 5.1 Setup
Models. Twelve neural networks, supported across four domains (graphs, hypergraphs, simplicial complexes, and cell complexes), are benchmarked. These include three GNNs (GCN, GIN, and GAT), three hypergraph neural networks (EDGNN, AllSetTransformer, and UniGNN2), three simplicial neural networks (SCN, SCCN, and SCCNN), and three cell complex neural networks (CCXN, CWN, and CCCN). Details on these architectures and their hyperparameters appear in Appendix C, while the number of learnable parameters for each best model configuration can be found in Table 7.
@@ 5.2 Main Results
As seen from Table 1, higher-order neural networks (based on hypergraphs, simplicial, and cell complexes) achieve the best performance on fifteen of twenty-two datasets, whereas GNNs achieve the best performance on six datasets, and tie on the Unemployment dataset. GNNs perform best on node regression in the majority of cases (five out of seven). These best results obtained by GNNs are closely matched by TNNs, since the latter achieve metrics within one standard deviation from the former. In contrast, in nine out of sixteen datasets TNNs outperform GNNs, and attain performance metrics that are higher by more than one standard deviation with respect to GNNs. In other words, in situations where higher-order networks outperform GNNs, the performance gap is more pronounced. It is also noted that, for demonstration purposes, only one fixed lifting is considered to transform graph data to each of the considered topological domains (see Appendix C.4). These results suggest that TNNs have an advantage over GNNs, even without lifting optimization. More importantly, for the context of this paper, the benchmarks demonstrate the degree of comparisons that can be performed with TopoBench across models and datasets.
@@ 5.3 Ablation Study
This ablation study examines how different readout strategies influence performance in neural networks built on higher-order domains, highlighting the importance of node-level signal updates and pooling choices. First, graph and hypergraph (neural network) models differ from simplicial and cell complex (neural network) models in terms of the domains and, subsequently, representations they support. Graph and hypergraph models can output two types of representations: node representations and edge or hyperedge representations. In contrast, the output of simplicial and cell models depends on the different types of cells present (00-cell up to $n$-cells) and on the model itself. For example, a simplicial or cell complex model may process an $n$-cell input but may not produce an $n$-cell output. The backbone_wrapper in TopoBench addresses these differences in the underlying domains of the models.
@@
Nevertheless, to fairly compare different neural network architectures, this second difference must be addressed. To that end, this ablation study considers two types of readouts to enable a rigorous evaluation: direct readout (DR), where the downstream task is performed directly over the 00-cell representation, and signal down-propagation (SDP), where information from higher-order cell representations is iteratively fused down to 00-cell representations using appropriate incidence matrices, followed by a linear projection over the concatenated $(n-1)$-cell signal and the fused $(n-1)$-cell representation. For instance, if a simplicial or cell complex model outputs 00-cell, $1$-cell, and $2$-cell representations, the signal propagates from $2$-cells to $1$-cells and then from $1$-cells to 00-cells during readout. The downstream task is then performed over the updated 00-cell representations. [^4] See Appendix A.3 for an introduction to TNNs and Higher-Order Message Passing on topological domains.
Table 2 shows that the best-performing readout type depends on how a model propagates signals internally. For example, the CWN model does not update 00-cell representations, so the SDP strategy performs notably better. Conversely, CCCN, SCCNN, and SCN propagate information to 00-cells, making SDP readout yield only small or negligible changes in performance. Further details are available in Appendix C.4.
@@
Table 2: An ablation study compares the performance of CWN, CCCN, SCCNN, and SCN models on various datasets using two readout strategies, direct readout (DR) and signal down-propagation (SDP). SDP generally enhances CWN performance, whereas the effect of SDP on CCCN, SCCNN, and SCN varies based on their internal signal propagation mechanisms. Means and standard deviations of performance metrics are shown. The best results are shown in bold for each model and readout type.
@@ 6 Concluding Remarks, Limitations, and Future Work
While TopoBench already addresses several challenges in TDL, it also has limitations that point to promising directions for future enhancements. One area is the integration of learnable liftings, which are currently not implemented. This direction could enable task-specific topological representations learned dynamically from data. A second limitation lies in the broader scarcity of standardized, real-world higher-order datasets. Although TopoBench incorporates numerous datasets for hypergraph, simplicial, and cell complexes, this remains an active area of expansion. Providing more built-in higher-order datasets will further streamline research in TDL. Finally, extending the set of evaluation metrics beyond classification or regression accuracy to include more TDL-specific measures of expressivity, explainability, and fairness (Papamarkou et al., 2024) is another avenue of growth.
Moving forward, the modular design of TopoBench invites contributions from the community. Researchers and practitioners are encouraged to contribute to TopoBench by introducing new learnable liftings, adding datasets, and developing specialized performance metrics. These efforts will not only strengthen the benchmarking ecosystem of TDL but also help drive innovation in topological deep learning more broadly.
@@
All aspects of library installation and development are described in the README.md file. To replicate the experiments reported in this paper, refer to the ‘Experiments Reproducibility’ section in README.md. Additional tutorials in the ‘Tutorials’ section illustrate how to integrate new models, datasets, and transforms within TopoBench.

### v3

@@ 1 Introduction
In geometric deep learning (GDL; Bronstein et al., 2021), graph neural networks (GNNs; Zhou et al., 2020) have demonstrated impressive capabilities in processing relational data represented as graphs. However, because graphs represent relationships through edges, they inherently capture only pairwise interactions, which can be a limiting factor. For example, social interactions often involve groups of individuals rather than just pairs, and electrostatic interactions in proteins can span multiple atoms. Topological deep learning (TDL; Papamarkou et al., 2024; Bodnar, 2023; Hajij et al., 2023b; Papillon et al., 2023) offers a framework for modeling complex systems characterized by such multi-way relations among components, leveraging to that end higher-order discrete topological domains (such as simplicial and cell complexes, see Section 2). Topological neural networks (TNNs; Feng et al., 2019; Bunch et al., 2020; Hajij et al., 2020; Bodnar et al., 2021a; Ebli et al., 2020; Schaub et al., 2021; Bodnar et al., 2021b; Chien et al., 2021), which are part of TDL, have found applications in numerous fields that involve higher-order relational data such as social networks (Knoke and Yang, 2019), protein biology (Jha et al., 2022), physics (Wei and Fink, 2024), and computer networks (Bernárdez et al., 2025). TNNs have also shown their potential in various machine learning tasks (Dong et al., 2020; Barbarossa and Sardellitti, 2020; Chen et al., 2022; Roddenberry et al., 2021; Telyatnikov et al., 2025; Giusti et al., 2023).
@@ 2 Background
Figure 2: Topological Deep Learning Domains. Nodes in blue, (hyper)edges in pink, and faces in dark red. Figure adapted from Papillon et al. (2023).
@@ 3 Existing Software
Various tools also exist for higher-order domains. For hypergraphs, simplicial complexes, and other topological structures, HyperNetX (Liu et al., 2021), XGI (Landry et al., 2023), DHG (Feng et al., 2019), and TopoX (Hajij et al., 2024) each focus on different facets. HyperNetX facilitates hypergraph computations, whereas XGI supports both hypergraphs and simplicial complexes. DHG implements deep learning algorithms for graphs and hypergraphs. TopoX is a suite of three packages—TopoNetX, TopoEmbedX, and TopoModelX—that provide broader support for hypergraphs, simplicial, cellular, path, and combinatorial complexes (Hajij et al., 2023a). TopoNetX facilitates constructing and computing on these domains, including working with nodes, edges, and higher-order cells; TopoEmbedX embeds higher-order domains into Euclidean spaces, while TopoModelX implements most TNNs surveyed in Papillon et al. (2023).
@@
TopoBench Contextualization
TopoBench leverages and extends this existing software ecosystem to provide a unified benchmarking infrastructure for TDL. The framework directly integrates established libraries including NetworkX for graph computations and the TopoX suite—TopoNetX for higher-order structure construction and TopoModelX for TNN implementations. TopoBench also incorporates graph-based models from PyG and enables seamless integration of models from original research repositories, providing unprecedented flexibility for TDL evaluation.
While these existing packages provide essential building blocks, TopoBench introduces novel capabilities that address critical gaps in the TDL software ecosystem. Unlike OGB’s focus on graph learning, TopoBench provides comprehensive data management for topological domains, including automated dataset downloading, storage, and processing capabilities. The framework introduces automated lifting transformations that extend beyond TopoNetX’s manual construction capabilities, enabling seamless data connectivity transformations between topological domains with integrated feature handling. Additionally, TopoBench offers unified mini-batching across all topological structures through a shared dataloader and streamlined configuration systems for experiment setup—capabilities absent from current TDL software.
These innovations collectively establish TopoBench as the first comprehensive benchmarking framework for TDL. The framework’s unified data representation enables consistent treatment of diverse topological structures, allowing researchers to evaluate models across different domains using standardized procedures. This approach transforms the fragmented TDL software landscape into a cohesive research environment, providing the reproducible benchmarking infrastructure that the rapidly evolving field requires.
@@ 4 The TopoBench Library: Module Outline, Datasets and Liftings
TopoBench implements a unified and flexible workflow that facilitates the addition of new datasets, data manipulation and preprocessing methods (collectively referred to as transforms), deep learning models, as well as custom metrics and losses. This design ensures applicability across a wide range of tasks and enables a broad cross-domain comparison, currently lacking in the TDL literature. Each module within TopoBench is assigned a distinct role while maintaining a consistent input-output structure, which provides a modular interface across all topological domains. Figure 1 outlines the TopoBench modules, grouped by functionality into data, model, training, and communication components. Algorithm 1 illustrates the TopoBench execution pipeline in pseudo-code.
@@
6:   $n$.Encoder(cfg.model),
7:   $n$.Backbone(cfg.model),
8:   $n$.BackboneWrapper(cfg.model),
9:   $n$.Readout(cfg.model),
@@ 4.1 TopoBench Modules
The readout module converts latent representations from the neural network into final predictions. The Loss module defines a loss function (from the PyTorch library, or customized), while the Optimizer module configures the optimizer and scheduler. This design allows seamless use of any optimizer and scheduler from torch.optim, thereby supporting flexible and robust training. Finally, the evaluator module, built upon torchmetrics, provides metrics for both classification and regression tasks –while also allowing for tailored ones for specific datasets and tasks. Notably, the flexibility of these modules enable researchers to implement topology-specific evaluation criteria as needed for their particular applications.
@@ 4.2 Datasets
Datasets with higher-order interactions. Several datasets with higher-order interactions are included in TopoBench, showcasing its capabilities to handle data supported on hypergraphs, simplicial complexes, and other topological domains. The MANTRA dataset (Ballester et al., 2024) is part of TopoBench, offering over 43,138 two-dimensional and 249,000 three-dimensional triangulations of surfaces and manifolds, which can be used, for example, as features on a simplicial complex. In addition, the widely used AllSet hypergraph datasets (Chien et al., 2021)—Cora-Cocitation, Citeseer-Cocitation, PubMed-Cocitation, Cora-Coauthorship, and DBLP-Coauthorship—are integrated, following the same preprocessing as HyperGCN (Yadati et al., 2019). These hypergraphs group documents co-authored or co-cited together into single hyperedges. Collectively, these examples illustrate how TopoBench supports data beyond traditional graph pairwise interactions.
@@ 4.3 Topological Liftings
Table LABEL:tab:submissions in Appendix B provides a comprehensive list of all the liftings currently implemented in TopoBench. Currently, TopoBench supports 11 structural liftings targeting simplicial complexes; 2 targeting cell complexes; 10 moving to hypergraphs; and 3 from a domain to combinatorial complexes. Notably, TopoBench’s modular design simplifies the integration of additional liftings, ensuring the framework’s adaptability to evolving research needs.
@@ 5 Numerical Experiments
This section presents numerical experiments that illustrate the breadth of TopoBench’s functionality by performing a cross-domain comparison. The overall setup is first described, then results from benchmarking various graph, hypergraph, and TNNs are reported, and an ablation study on signal propagation is presented to demonstrate how TopoBench supports comparisons in TDL.
@@ 5.1 Setup
Models. Twelve neural networks, supported across four domains (graphs, hypergraphs, simplicial complexes, and cell complexes), are benchmarked. These include three GNNs (GCN, GIN, and GAT), three hypergraph neural networks (EDGNN, AllSetTransformer, and UniGNN2), three simplicial neural networks (SCN, SCCN, and SCCNN), and three cell complex neural networks (CCXN, CWN, and CCCN). Details on these architectures and their hyperparameters appear in Appendix C. In particular, the number of learnable parameters for each best model configuration can be found in Table 8, while the corresponding runtimes are provided in Table 9.
@@ 5.2 Main Results
As seen from Table 1, higher-order neural networks (based on hypergraphs, simplicial, and cell complexes) achieve the best performance on fifteen of twenty-two datasets, whereas GNNs achieve the best performance on six datasets, and tie on the Unemployment dataset. GNNs perform best on node regression in the majority of cases (five out of seven). These best results obtained by GNNs are closely matched by TNNs, since the latter achieve metrics within one standard deviation from the former. In contrast, in nine out of sixteen datasets TNNs outperform GNNs, and attain performance metrics that are higher by more than one standard deviation with respect to GNNs. In other words, in situations where higher-order networks outperform GNNs, the performance gap is more pronounced. It is also noted that, for demonstration purposes, only one fixed lifting is considered to transform graph data to each of the considered topological domains (see Appendix C.4). These results suggest that, even without lifting optimization, TNNs have an advantage over GNNs in terms of performance, although it is worth emphasizing that overall they also tend to be less efficient in terms of memory usage and computational time than graph-based counterparts (see Appendix C.4 for a more detailed analysis). However, and more importantly for the context of this paper, the benchmarks demonstrate the degree of comparisons that can be performed with TopoBench across models and datasets.
@@
Remark.
Notably, OOM results are originated when lifting large, densely connected graphs to higher-order domains, showcasing the scalability issues of the liftings leveraged in this analysis (i.e., clique and cycle liftings to simplicial and cellular domains, respectively).
@@ 5.3 Ablation Study
This ablation study examines how different readout strategies influence performance in neural networks built on higher-order domains, highlighting the importance of node-level signal updates and pooling choices. First, graph and hypergraph (neural network) models differ from simplicial and cell complex (neural network) models in terms of the domains and, subsequently, representations they support. Graph and hypergraph models can output two types of representations: node representations and edge or hyperedge representations. In contrast, the output of simplicial and cell models depends on the different types of cells present (0-cell up to $n$-cells) and on the model itself. For example, a simplicial or cell complex model may process an $n$-cell input but may not produce an $n$-cell output. The backbone_wrapper in TopoBench addresses these differences in the underlying domains of the models.
@@
Nevertheless, to fairly compare different neural network architectures, this second difference must be addressed. To that end, this ablation study considers two types of readouts to enable a rigorous evaluation: direct readout (DR), where the downstream task is performed directly over the 0-cell representation, and signal down-propagation (SDP), where information from higher-order cell representations is iteratively fused down to 0-cell representations using appropriate incidence matrices, followed by a linear projection over the concatenated (n1)(n-1)-cell signal and the fused (n1)(n-1)-cell representation. For instance, if a simplicial or cell complex model outputs 0-cell, 11-cell, and $2$-cell representations, the signal propagates from $2$-cells to $1$-cells and then from $1$-cells to 0-cells during readout. The downstream task is then performed over the updated 0-cell representations. [^3] See Appendix A.3 for an introduction to TNNs and Higher-Order Message Passing on topological domains.
Table 2 shows that the best-performing readout type depends on how a model propagates signals internally. For example, the CWN model does not update 0-cell representations, so the SDP strategy performs notably better. Conversely, CCCN, SCCNN, and SCN propagate information to 0-cells, making SDP readout yield only small or negligible changes in performance. Further details are available in Appendix C.4.
@@
Table 2: This ablation study compares the performance of CWN, CCCN, SCCNN, and SCN models on various datasets using two readout strategies, direct readout (DR) and signal down-propagation (SDP). SDP generally enhances CWN performance, whereas the effect of SDP on CCCN, SCCNN, and SCN varies based on their internal signal propagation mechanisms. Means and standard deviations of performance metrics are shown. The best results are shown in bold for each model and readout type.
@@
5.4 Higher-Order Datasets
Appendix D presents additional illustrative experiments conducted on 13 datasets included in TopoBench, spanning a broad range of hypergraph datasets (for classification tasks) and simplicial datasets (for both classification and regression tasks). The evaluation protocol follows the setup described in Section 5.1, with the exception of structural and feature liftings, as these datasets natively possess higher-order topologies and include features on higher-order cells.
For the hypergraph datasets, no single model consistently outperforms others across all benchmarks. Among the evaluated models, AllSetTransformer achieves the best performance on 5 out of 10 datasets. For the simplicial MANTRA family of datasets, the results demonstrate that topological tasks are more effectively modeled by TNNs, whereas standard GNN baselines fail to capture the intricate topological structures, resulting in lower performance on purely topological tasks.
@@ 6 Concluding Remarks, Limitations, and Future Work
While TopoBench already addresses several challenges in TDL, it also has limitations that point to promising directions for future enhancements. One area is the implementation of learnable liftings, which are supported by  TopoBench. This direction could enable task-specific topological representations learned dynamically from data. A second limitation lies in the broader scarcity of standardized, real-world higher-order datasets. Although TopoBench incorporates numerous datasets for hypergraph, simplicial, and cell complexes, this remains an active area of expansion. Providing more built-in higher-order datasets will further streamline research in TDL.
Another potential direction is to perform an exhaustive exploration of optimal liftings per combination of domains, datasets, and models. In fact, OOM values showcase the scalability limitations of the most common used strategies to lift graphs into simplicial and cellular domains (i.e., clique and cycle liftings, respectively). Finally, extending the set of evaluation metrics beyond classification or regression accuracy to include more TDL-specific measures of expressivity, explainability, and fairness (Papamarkou et al., 2024) is another avenue of growth –and these modules have been designed to be easily extendable.
Moving forward, the modular design of TopoBench invites contributions from the community. Researchers and practitioners are encouraged to contribute to TopoBench by introducing new learnable liftings, adding datasets, and developing specialized performance metrics. Moreover, to mitigate the aforementioned scalability issues, several strategies can be explored –e.g. pruning the input graphs prior to lifting, employing scalable lifting mechanisms—such as those explored in the ICML 2024 TDL Challenge (Bernárdez et al., 2024)—and applying mini-batching techniques to higher-order structures in transductive settings (analogous to those used in GNN modeling). These efforts will not only strengthen the benchmarking ecosystem of TDL but also help drive innovation in topological deep learning more broadly –as already shown in the recents works of TopoTune (Papillon et al., 2025) and HOPSE (Carrasco et al., 2025), both of which leverage TopoBench framework to push the boundaries of TDL.
@@
All aspects of library installation and development are described in the README.md file. To replicate the experiments reported in this paper, refer to the ‘Experiments Reproducibility’ section in README.md. Additional tutorials in the ‘Tutorials’ section illustrate how to integrate new models, datasets, learnable liftings, and transforms within TopoBench.
@@
Author Contributions
L. Telyatnikov and G. Bernárdez contributed equally to this work as the main authors and lead developers. The conceptualization of the TopoBench project was a collaborative effort by L. Telyatnikov, G. Bernárdez, M. Montagna, N. Miolane, T. Papamarkou, M. Hajij, G. Zamzmi, M. T. Schaub, and S. Scardapane. The core development and implementation of the benchmark were carried out by L. Telyatnikov, G. Bernárdez, M. Montagna, M. Carrasco, P. Vasylenko, M. Papillon, and N. Miolane. The experiments were led by L. Telyatnikov with support from G. Bernárdez. The manuscript was written by T. Papamarkou, L. Telyatnikov, and G. Bernárdez, with significant writing contributions to various sections from S. Scardapane, M. Hajij, G. Zamzmi, and M. T. Schaub. All other authors contributed to the TopoBench ecosystem through their winning submissions (i.e. lifting implementations) to the ICML TDL Challenge 2024 (Bernárdez et al., 2024).
Acknowledgments and Disclosure of Funding
M. Papillon, G. Bernárdez and N. Miolane acknowledge support from the National Science Foundation, Award DMS-2134241. M. Papillon and N. Miolane acknowledge funding from the National Science Foundation, Award DMS-2240158 and from the Noyce Foundation. M. Papillon acknowledges the support of the Natural Sciences and Engineering Research Council of Canada. M. Hajij acknowledges support from the National Science Foundation, award DMS-2134231. M. T. Schaub acknowledges funding by the European Union (ERC, HIGH-HOPeS, 101039827). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.